OpenStudy (anonymous):

Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

OpenStudy (anonymous):

2x + y = 10 multibly by -4 3x + 4y = 25 -8x-4y=-40 3x +4y=25 add the two equations -5x = -15 x=3 sub this in first equation 2× 3+y=10 y = 4